Twice Andy's age plus 3 is 25. How old is Andy?
Answer by umail08(87)

x = Andy’s age
2x = twice Andy’s age
2x + 3 = twice Andy’s age plus 3
2x + 3 = 25 … equation to solve
2x + 3 – 3 = 25 – 3
2x = 22
2x/2 = 22/2
x = 11 … Andy’s age